OpenStudy (anonymous):

Find the length of the hypotenuse with the given leg lengths. Round to the nearest tenth. a=3,b=5 a=11,b=14

OpenStudy (anonymous):

first answer is 5.8 second answer is 17.8

OpenStudy (anonymous):

use pythagoris theorem

OpenStudy (anonymous):

\[\sqrt{3^2+5^2}=5.8\]

OpenStudy (anonymous):

\[\sqrt{11^2+14^2}\]=17.8